A perfectly competitive firm has the cost function TC = 1000 +  2Q  + 0.1  Q2. What is the lowest price at which this firm can break even? 


1
Expert's answer
2021-05-19T20:29:23-0400

"MC = 2 + 0.2Q"


"AC = \frac {1000} {Q} + 2 + 0.1Q"


Set MC = AC.


For minimum AC, or "2 + 0.2Q = \frac {1000} {Q} + 2 + 0.1Q, or Q = 100" , and at that point, AC = $22.


This is the lower price that the firm can break even.